OpenStudy (anonymous):

A quadratic equation is shown below: x2 - 8x + 13 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? Subtract 5 from both sides of the equation Subtract 3 from both sides of the equation Add 5 to both sides of the equation <--my answer Add 3 to both sides of the equation

OpenStudy (radar):

This looks like "completing the sum of square" problem, or at least some training in getting ready to use the procedure. Since the x^2 term already has a 1 for its coefficient you can start like this. Take the coefficient of the x term (-8), and take half of it, that would be a -4, now square that giving you a 16. that is the key. Now you have already a 13, you need a 16. What would you do to get a 16. This would result is x^2 -8x + 16 = 3 giving you (x-4)^2 = 3 which is in the desired form (x-p)^2 = q.

OpenStudy (anonymous):

so my answer was correct ?

OpenStudy (radar):

No you added 5 giving you an 18 not a 16.

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

add 3

OpenStudy (radar):

You want to add just the amount to make a perfect square. A 3 works nicely, noticed since you added a 3 to the left side of the equal sign, you have to add a 3 to the right of the equal sign to keep the equality balanced.

OpenStudy (anonymous):

so D is correct

OpenStudy (radar):

If D is adding a 3, as I am trying to explain, then it is correct.